OpenStudy (anonymous):

solve the given equation |2x+7|=1

OpenStudy (anonymous):

when |2x+7| is positive, +(2x+7)=1 2x=1-7 2x=-6 x=-3 and, when |2x+7| is negative, -(2x+7)=1 -2x-7=1 -2x=1+7 -2x=8 x=-4

OpenStudy (anonymous):

the two are the answers because of the | | sign?

OpenStudy (anonymous):

yes u have to consider two situations, when modulus values is positive and when dis va;lue is negative.i hope dis would be the answer.

OpenStudy (anonymous):

thanks :)